OpenStudy (abbycross167):

Can someone please help me? I really need help... I will fan and Medal Question~ Simplify the following exponential expressions. 5x^6y^4 --------- -20xy^-2z

OpenStudy (abbycross167):

@phi would you mind helping me?

OpenStudy (phi):

let's hope OS behaves. meanwhile, what is 5/-20 ? (remember this is plus divided by minus, so the answer is going to be negative)

OpenStudy (abbycross167):

-1/4

OpenStudy (abbycross167):

0?

OpenStudy (phi):

oh! I meant anything divided by itself is 1!

OpenStudy (abbycross167):

oh okay, haha that's what I thought it was supposed to be! ok so x/x=1

OpenStudy (phi):

all of these problems use that idea (but there are short-cuts) if you see \[ \frac{x}{x}\] that is 1 if you see \[ \frac{x\cdot x \cdot x}{x} \] that is the same as \[ \frac{x}{x} \cdot x\cdot x \] and x/x is 1, so you get \[ 1\cdot x \cdot x\] and we can ignore multiply by 1 (it does not "do anything") in other words you get \[ x \cdot x\] does that make sense?

OpenStudy (phi):

you have in your problem \[ \frac{x^6}{x} \] the "little 6" means x times itself 6 times so (if we wanted) we could write this as \[ \frac{x \cdot x \cdot x \cdot x \cdot x \cdot x}{x}\] and the x in the bottom will "cancel out" one of the x's up top that will leave only 5 x's up top we would write that in "short-hand" as \[ x^5\]

OpenStudy (phi):

next we do the y's \[ \frac{y^4}{y^{-2}} \] the bottom y is y^-2 the negative number -2 is a bit tricky. You have to learn this rule: if you move a number or letter from bottom to top, you change the sign of the exponent in other words, to make that -2 turn into a +2, we "move" the y^-2 up top \[ \frac{y^4 y^2}{1} \] that is 4 y's multiplied together times 2 more y's so y*y*y*y * y*y how many y's is that altogether ?

OpenStudy (abbycross167):

y=6

OpenStudy (phi):

yes, but not y=6 you have y*y*y*y*y*y and if you don't like to write that out, just write y and put an ^ and a 6 next to it y^6 (or if you use the equation editor, or write it) make the 6 small in the upper right. \[ y^6\]

OpenStudy (phi):

do you understand how the "short-cut" way to write it works ?

OpenStudy (abbycross167):

yes sir

OpenStudy (phi):

do you know how to write x*x*x*x using the short way ?

OpenStudy (abbycross167):

x^4?

OpenStudy (phi):

yes. and do you know how to write y^3 the "long way" ?

OpenStudy (abbycross167):

y*y*y do you mean this way?

OpenStudy (phi):

yes. you are doing good. now can you simplify \[ \frac{x\cdot x\cdot x}{x\cdot x\cdot x} \] notice you can "cancel" each "pair of x" (one in the top with one in the bottom)

OpenStudy (abbycross167):

yes sir

OpenStudy (phi):

what does it simplify to ?

OpenStudy (abbycross167):

@phi it equals 0?

OpenStudy (phi):

\[ \frac{x\cdot x\cdot x}{x\cdot x\cdot x} \\= \frac{x}{x}\cdot\frac{x}{x}\cdot\frac{x}{x} \\ = 1\cdot 1\cdot 1\\=1 \]